OpenStudy (anonymous):

solve the following using addition: x=y+1400 0.07x+0.04y=670

OpenStudy (angela210793):

x=y+1400 x-y=1400(multiply with 0.04) 0.04x-0.04y=56 0.07x+0.04y=670---> 0.07x+0.04y=670 0.07x+0.04y=670 add 0.11x=726 x=726/0.11=6600 y=6600-1400=5200

OpenStudy (anonymous):

thank you!

OpenStudy (angela210793):

uw

OpenStudy (anonymous):

Good Job Angela!!! U deserve a medal !!!!

OpenStudy (angela210793):

^_^ thnx :)

OpenStudy (radar):

She gets a medal.

OpenStudy (angela210793):

I got 2 :)

OpenStudy (anonymous):

I hope rhoffman gives u one too!!

OpenStudy (anonymous):

how do i mark it as a good answer? new to this...

OpenStudy (anonymous):

nm - browser issue - button wasn't there

OpenStudy (angela210793):

hmm thnx ^_^